Determine all of the points on the curve

y = 3x2 + 24x - 3 where the slope of the tangent line is 6


1
Expert's answer
2021-02-28T07:43:39-0500

Consider the function "y=3x^2+24x-3"


Differentiate the function "y" with respect to "x" as,


"\frac{dy}{dx}=\frac{d}{dx}(3x^2+24x-3)"


"=3(2x)+24(1)"


"=6x+24"


Since, the slope of the tangent line is "6" , therefore,


"\frac{dy}{dx}=6"


"6x+24=6"


"6x+24-24=6-24"


"6x=-18"


"x=-3"


At "x=-3" , "y=3(-3)^2+24(-3)-3=-48"


Therefore, the point on the curve in ordered pair is "(-3,-48)".